# “I Want It That Way”: Enabling Interactive Decision Support Using Large Language Models and Constraint Programming

### v1

@@ 2.3. LLMs and Optimization
Large Language Models (LLMs) have emerged as a promising tool for helping accomplish a number of tasks such as generating code (Chen et al., 2021), writing (Yuan et al., 2022), and visualizing data (Dibia, 2023). However, they are well known to suffer from hallucinations, seemingly plausible text predictions that are in fact non-factual (Welleck et al., 2019), and poor mathematical reasoning ability (Qian et al., 2022). Recent work have aimed to address these shortcomings by helping LLMs decompose complex tasks into simpler steps (e.g., (Wei et al., 2022; Yao et al., 2022, 2023)) and augmenting them with external tools (e.g., (Schick et al., 2023; Thoppilan et al., 2022; Nakano et al., 2021)) such as a code interpreter (see Mialon et al. (2023) for a survey of other related approaches). Our work builds upon this line of work by decomposing the complex problem of suggesting meeting times into structured sub-components (detailed in Section 4). We also employ an optimization solver as an external tool to perform complex reasoning over meeting times. This is in contrast to other existing working on collaborative decision making between LLMs and humans such as Lin et al. (2023) which requires the LLM to generate the solutions.
@@
Mathematical optimization is a powerful tool used in various industrial applications (Bertsimas and Tsitsiklis, 1997). In this paper, our focus is on (Weighted) Constraint Programming (Rossi, 1999), a modeling paradigm where the optimization problem involves a set of arbitrary constraints with different weights. The objective is to find a solution with the optimal score, determined by the total weight of satisfied constraints. Traditionally, constraint programming demands a complete specification of the optimization model, including all constraints and their coefficients. This poses challenges for preference elicitation, where users’ constraints and preferences are often unknown and expressed within the context of proposed solutions. Recent research has addressed this by tackling constraint programming problems with incomplete information, requiring the elicitation of certain data and preferences (Tabakhi, 2017; Tabakhi et al., 2017; Tabakhi, 2021; Tabakhi et al., 2022; Xiao, 2020). However, existing approaches primarily focus on the algorithmic aspects of what information to elicit, rather than the mechanism for eliciting it. This distinction is crucial, particularly in settings like meeting scheduling, where non-expert users may lack the modeling expertise to formulate the constraints they want in the model.
@@
Natural language processing (NLP) holds promise in simplifying interactions with optimization models for non-expert users. Recent work has employed LLMs for users to query models in areas like supply chain management (Li et al., 2023) and identifying infeasible subsystems (Chen et al., 2023). However, these chat-based systems primarily aid in understanding existing models, not constructing new ones. To address the challenge of formulating optimization problems, recent research focuses on translating natural language descriptions into formal optimization models. Ramamonjison et al. (2022)) initiated this effort, leading to the NL4Opt competition at NeurIPS and related work on named entity recognition and semantic parsing (Ning et al., 2023; Tsouros et al., 2023; Dakle et al., 2023; Prasath and Karande, 2023). These approaches typically address static contexts, translating a single text paragraph into an optimization model. In contrast, our study focuses on an interactive model, capturing user preferences during optimization model construction. We emphasize translating natural language into Python constraints within a constraint programming framework, distinct from traditional named entity recognition for linear and integer programming.
@@ 3. Preference Construction: Understanding Contextual Scheduling Preferences
We first investigated preference construction in the context of meeting scheduling. Specifically, we aimed to identify and characterize the types of contextualized preferences elicited in response to given time suggestions. In the absence of an existing interactive system through which we could analyze users preferences in response to dynamic time suggestions, we conducted a diary study at a large technology company that elicited user preferences in response to static time suggestions from an existing calendar tool. The findings of our diary study were used to inform the design of the system (Section 4) and create a new quantitative benchmark for its performance (Section 5). We also evaluate the system in the context of dynamic time suggestions in the user study outlined in Section 6.
@@ 3.1. Study Protocol
To capture contextual scheduling preferences for a given meeting, participants were asked to complete a diary entry for each meeting they scheduled during a one-week study period. Participants were instructed to first phrase their initial meeting scheduling request to a pretend chat-based interface (e.g., “I want to schedule a 30 minute 1:1 with my manager within the next week”), and then enter the needed information into a widely used commercial calendaring tool that generates time suggestions for a meeting. In response to the suggested times, participants were then asked to express any preferences up to 5 times (e.g., “I can’t meet before 10am so that I can drop my kids off at work”). They were also asked to reflect on whether there was information missing from their calendar that could have captured their stated preference. Finally, participants were also asked to include information about the meeting they were scheduling including the number of attendees, its duration, and the time horizon in which it was to be scheduled.
We recruited 64 participants through an e-mail advertisement at a large technology company who scheduled their own meetings (i.e. did not have an administrative assistant) and had at least 2 meetings per week. Participants self-reported as working in the roles of a manager (39%), program manager (30%), developer (17%), and researcher (9%). 39% of participants reported attending under 20 meetings per week, 37% attended between 20-30, and the remainder reported over 30 meetings per week. We collected diary entries for 114 unique meetings with averages of 3.95 attendees and 35 minutes. 58% of the meetings were expected to be fully virtual, 20% hybrid, and 4% fully in-person. In total, we collected 114 initial phrasing of the scheduling request, 211 contextual preference expressions, and 197 missing information descriptions. We donated \$300 to charity for participation in the study. The study design and protocol were approved by the leading institution’s Institutional Review Board (IRB).
@@ 3.2. Analysis & Findings
We qualitatively analyzed all preference and missing information entries (Charmaz, 2006). Researchers conducted open-coding of 30% of the entries. Initial codes were iterated and grouped into coarser categories of preferences and missing information to develop a codebook which was subsequently used to conduct closed-coding of the remainder of the data. While our survey asked users to express preferences in response to suggested times, we also observed users requesting additional actions such as querying calendar data (e.g., “Are [external participants] available during this time?”), drafting an agenda (e.g., “Draft the agenda based on AIs from previous meeting with the person”), or sending messages (e.g., “send a message asking to move the meeting earlier rather than later”). We restricted our analysis to focus only on preferences, as opposed to these auxiliary actions. We also labeled each of the initial phrasings of the scheduling requests for the presence of the duration of the meeting, the attendees, and the expected time frame. We then reviewed the labeled and coded entries to highlight four key takeaways from our analysis.
@@
3.2.1. Scheduling preferences are diverse.
@@
3.2.2. Scheduling requests are vague.
@@
  • Personal Commitment: Scheduled events or obligations that were missing from the user’s work calendar (e.g., “I was planning to be oof but didn’t make it in the calendar (P70)”).
@@
Figure 1(b) shows a breakdown of the missing information expressed during the diary study along these axes. While an ideal calendar system may be able to internalize some of this missing information (i.e. by asking users to input all their personal commitments), the scope and diversity of the missing information makes capturing it all in existing systems near impossible. This highlights the importance of a flexible interface through which users can inform the scheduling system of context and information missing from a user’s calendar.
3.2.4. Users value temporal diversity.
@@ 4. Preference Incorporation: MeetMate System
Towards a flexible, contextual, and interactive decision support system, we introduce MeetMate, a chat-based meeting scheduling tool built in Python that allows users to interactively refine meeting time suggestions via natural language requests. A key component of our approach is an underlying constraint programming optimization model to generate new time suggestions (detailed in Section 4.3), which performs the complex mathematical reasoning required to find suitable times. We model each scheduling preference as a function, called a constraint, that checks whether a candidate meeting time meets the preference (i.e., “Not scheduled before 2pm”). Each constraint also has an associated weight. Intuitively, higher weights correspond to constraints that are more important. Instead of requiring users to specify an exact numerical weight for each preference, the system instead keeps an underlying prioritized list of the constraints which are then translated into weights such that constraints with higher priority are always satisfied before those of lower priority. Each constraint is formalized as a Python function that is used by a solver to generate new time suggestions. To embed natural language scheduling preferences into Python functions, we leverage LLMs to translate user chat messages to operations on the weighted list of scheduling constraints. This hybrid framework capitalizes on the benefits of both LLMs and Optimization, using LLMs to flexibly embed naturalistic scheduling constraints into Python functions and using optimization to generate new time suggestions that are responsive to the embedded preferences. This occurs within an interactive chat environment which allows the system to be both contextualized to a given scheduling instance and interactive.
@@
When a user goes to schedule a meeting, they must first enter the list of attendees, the duration, and the time horizon for the meeting into a form. Using a form ensures that users include the minimum viable amount of information to schedule a meeting, and alleviates the vague scheduling behavior we discovered during the diary study. Once a user enters the information, the system generates an initial time suggestion that maximizes the availability of the desired attendees. The system returns the time suggestion and initiates a chat so that users can interact with the system to refine the suggested meeting time. For every new chat message entered in the chat, the Constraint Management Component (Section 4.1) translates the request into an action for the system to take. One such action is the addition of a new constraint (see Section 4.2), which requires translating a natural language preference into a Python function. These Python functions are then used within a constraint programming solver (Section 4.3) to generate new time suggestions. An overview of the system is included in Figure 5.
@@ 4.1. Constraint Management
  • Add Constraint: This action generates a new scheduling constraint and calls the constraint generation functionality outlined in Section 4.2.
@@ 4.2. Constraint Generation
When the system decides to add a new scheduling constraint, the system calls two LLM-powered components. The first, dubbed Information Checker, checks whether the given scheduling constraint can be handled by the system. Given the diversity of user scheduling preferences and the wide range of external information needed to integrate them, it is unlikely that any system will be able to handle all scheduling constraints. The Information Checker acts as a safeguard to check whether the system has sufficient information to be able to handle the constraint. If it cannot, it identifies missing information that can either be queried from the user or states why the given constraint cannot be handled by the system. If the system can handle the constraint, it is then given to a Coder module that translates the scheduling constraint into Python code that can be used to check whether a candidate meeting time meets the constraint. Figure 6 shows a sample prompt for generating a new scheduling constraint.
Figure 5. Sample prompt for the constraint generation coder component. Orange highlight reflects scheduling instance-specific inputs. Green highlights the output of the LLM.

### v3

@@ 2.4. LLMs and Optimization
Large Language Models (LLMs) have emerged as a promising tool for helping accomplish a number of tasks such as generating code (Chen et al., 2021), writing (Yuan et al., 2022), and visualizing data (Dibia, 2023). LLMs have also been applied to convert human preferences directly into automated decision-making systems (Li et al., 2023b) or to solve constraint satisfaction problems (Yuksekgonul et al., 2024; Abdin et al., 2024). However, they are well known to suffer from hallucinations, seemingly plausible text predictions that are in fact non-factual (Yuksekgonul et al., 2024; Welleck et al., 2020), and poor mathematical reasoning ability (Qian et al., 2023). Recent work has aimed to address these shortcomings by helping LLMs decompose complex tasks into simpler steps (e.g., (Wei et al., 2022; Yao et al., 2023, 2024)) and augmenting them with external tools (e.g., (Schick et al., 2024; Thoppilan et al., 2022; Nakano et al., 2021)), such as a code interpreter. We refer to Mialon et al. (2023) for a survey of other related approaches. Our work builds upon this line of work by decomposing the complex problem of suggesting meeting times into structured sub-components (detailed in Section 4). We also employ an optimization solver as an external tool to perform complex reasoning over meeting times. This is in contrast to other existing work on collaborative decision-making between LLMs and humans, such as by Lin et al. (2024), which requires the LLM to generate the solutions.
@@
Mathematical optimization is a powerful tool used in various industrial applications (Bertsimas and Tsitsiklis, 1997). In this paper, our focus is on (Weighted) Constraint Programming (Rossi, 1999), a modeling paradigm where the optimization problem involves a set of constraints with different weights. The objective is to find a solution with the optimal score, determined by the total weight of satisfied constraints. Traditionally, constraint programming demands a complete specification of the optimization model, including all constraints and their coefficients. This poses challenges for preference elicitation, where users’ constraints and preferences are often unknown and expressed within the context of proposed solutions—as discussed earlier. Recent research has addressed this by tackling constraint programming problems with incomplete information, requiring the elicitation of certain data and preferences (Tabakhi, 2017; Tabakhi et al., 2017; Tabakhi, 2021; Tabakhi et al., 2022; Xiao, 2020). However, existing approaches primarily focus on the algorithmic aspects of what information to elicit, rather than the mechanism for eliciting it. This distinction is crucial, particularly in settings like meeting scheduling, where non-expert users may lack the modeling expertise to formulate the constraints they want in the model.
@@
Natural language processing holds promise in simplifying interactions with optimization models for non-expert users. Recent work has employed LLMs for users to query models in areas like supply chain management (Li et al., 2023a) and identifying infeasible subsystems (Chen et al., 2023). However, these chat-based systems primarily aid in understanding existing models, not constructing new ones. To address the challenge of formulating optimization problems, recent research has focused on translating natural language descriptions into formal optimization models. Ramamonjison et al. (2022) initiated this effort, leading to the NL4Opt competition at the NeurIPS conference and related work on named entity recognition and semantic parsing (Ning et al., 2023; Tsouros et al., 2023; Dakle et al., 2023; Prasath and Karande, 2023). These approaches typically address static contexts, translating a single text paragraph into an optimization model. In contrast, our study focuses on an interactive model, capturing user preferences during optimization model construction. We emphasize translating natural language into Python constraints within a constraint programming framework, distinct from traditional named entity recognition for linear and integer programming.
@@ 3. Preference Construction: Understanding Contextual Scheduling Preferences
We first investigated preference construction in the context of meeting scheduling. Specifically, we aimed to identify and characterize the types of contextualized preferences elicited in response to given time suggestions. In the absence of an existing interactive system through which we could analyze users preferences in response to dynamic time suggestions, we conducted a diary study at a large technology company that elicited user preferences in response to static time suggestions from an existing calendar tool. The findings of our diary study were used to inform the design of the system (Section 4) and create a new quantitative benchmark for its performance (Section 5). We evaluate the system in the context of dynamic time suggestions in the user study outlined in Section 6.
@@ 3.1. Study Protocol
To capture contextual scheduling preferences for a given meeting, participants were asked to complete a diary entry for each meeting they scheduled during a one-week study period. Participants were instructed to first phrase their initial meeting scheduling request to a pretend chat-based interface (e.g., “I want to schedule a 30 minute 1:1 with my manager within the next week”), and then enter the needed information into a widely used commercial calendaring tool that generates time suggestions for a meeting. In response to the suggested times, participants were then asked to express any preferences up to 5 times (e.g., “I can’t meet before 10am so that I can drop my kids off before work”). They were also asked to reflect on whether there was information missing from their calendar that could have captured their stated preference. Finally, participants were also asked to include information about the meeting they were scheduling including the number of attendees, its duration, and the time horizon in which it was to be scheduled. We recruited 64 adult participants based in the United States through an e-mail advertisement at a large technology company who scheduled their own meetings (i.e., did not have an administrative assistant) and had at least 2 meetings per week. Participants self-reported as working in the roles of a manager (39%), program manager (30%), developer (17%), and researcher (9%). 39% of participants reported attending under 20 meetings per week, 37% attended between 20-30, and the remainder reported over 30 meetings per week. We collected diary entries for 114 unique meetings with averages of 3.95 attendees and 35 minutes. 58% of the meetings were expected to be fully virtual, 20% hybrid, and 4% fully in-person (the modality of the remainder of meetings was uncertain at the time of scheduling). Despite no such requirement in the study instructions, all meetings scheduled during the diary study pertained to work scheduling. In total, we collected 114 instances of the initial phrasing of the scheduling request, 211 contextual preference expressions, and 197 missing information descriptions. We donated \$1 dollar to charity for each diary entry completed, totalling to \$114. Participants were informed of this reward scheme and the selected charity at the time of study enrollment. The study design and protocol were approved by the leading institution’s Institutional Review Board (IRB).
@@ 3.2. Analysis & Findings
We qualitatively analyzed all preference and missing information entries following  (Charmaz, 2006). Researchers conducted open-coding of 30% of the entries. Initial codes were iterated and grouped into coarser categories of preferences and missing information to develop a codebook which was subsequently used to conduct closed-coding of the remainder of the data. While our survey asked users to express preferences in response to suggested times, we also observed users requesting additional actions such as querying calendar data (e.g., “Are [external participants] available during this time?”), drafting an agenda (e.g., “Draft the agenda based on previous meeting with the person”), or sending messages (e.g., “send a message asking to move the meeting earlier rather than later”). We restricted our analysis to focus only on preferences, as opposed to these auxiliary actions. We also labeled each of the initial phrasings of the scheduling requests for the presence of the duration of the meeting, the attendees, and the expected time frame. We then reviewed the labeled and coded entries to highlight four key takeaways from our analysis.
@@
[…]
Two horizontal bar graphs, titled (a) Contextual Preferences and (b) Missing Information. Contextual Preferences shows scheduling preference category on the Y axis against percentage of responses on the X axis. Seven bars shown, ordered from largest to smallest category: Temporal- 54%, Existing Calendar- 16%, External Information- 8.5%, Relational- 8.5%, Attendance- 6%, Duration- 5%, and Facility- 2%. Missing Information shows missing information type on the Y axis against percentage of responses on the X axis. Seven bars shown, ordered from largest to smallest information type: Attendee Calendar- 22%, Personal Preference- 22%, Attendee Preference- 15%, Personal Commitment- 14%, Meeting Context- 12%, Timezone- 9%, and Other- 6%.
3.2.1. Scheduling preferences are diverse
@@
3.2.2. Scheduling requests are vague
@@ 3.2.3. Calendars are incomplete
  • Personal Commitment: Scheduled events or obligations that were missing from the user’s work calendar (e.g., “I was planning to be oof [^1] OOF refers to being ‘out of office‘. but didn’t make it in the calendar (P70)”).
@@
Figure 2(b) shows a breakdown of the missing information expressed during the diary study along these axes. While an ideal calendar system may be able to internalize some of this missing information (i.e., by asking users to input all their personal commitments), the scope and diversity of the missing information makes capturing it all in existing systems near impossible. This highlights the importance of a flexible interface through which users can inform the scheduling system of context and information missing from a user’s calendar.
3.2.4. Users value temporal diversity
@@ 4. Preference Incorporation: MeetMate System
Towards a flexible, contextual, and interactive decision support system, we introduce MeetMate, a chat-based meeting scheduling tool built in Python that allows users to interactively refine meeting time suggestions via natural language requests. A key component of our approach is an underlying constraint programming optimization model to generate new time suggestions (detailed in Section 4.3), which performs the complex mathematical reasoning required to find suitable times. We model each scheduling preference as a function, called a constraint, that checks whether a candidate meeting time meets the preference (e.g., “Not scheduled before 2pm”). Each constraint also has an associated weight. Intuitively, higher weights correspond to constraints that are more important. Instead of requiring users to specify an exact numerical weight for each preference, the system instead keeps an underlying ordered list of the constraints which are then translated into weights such that constraints with higher priority (i.e., earlier in the list) are always satisfied before those of lower priority (i.e., later in the list). Each constraint is formalized as a Python function that is used by a solver to generate new time suggestions. To embed natural language scheduling preferences into Python functions, we leverage LLMs to translate user chat messages to operations on the weighted list of scheduling constraints. This hybrid framework capitalizes on the benefits of both LLMs and Optimization, using LLMs to flexibly embed naturalistic scheduling constraints into Python functions and using optimization to generate new time suggestions that are responsive to the embedded preferences. This occurs within an interactive chat environment which allows the system to be both contextualized to a given scheduling instance and interactive.
@@
When a user begins to schedule a meeting, they must first enter the list of attendees, the duration, and the time horizon for the meeting into a form. Using a form ensures that users include the minimum viable amount of information to schedule a meeting, and alleviates the vague scheduling behavior we discovered during the diary study. This aligns with many existing calendar UIs which require users to specify a duration and attendees before generating candidate meeting times. Similar to popular institutional scheduling tools like Microsoft Outlook and Google Calendar, the system is given access to each attendee’s public free/busy times, but is not given information about the meetings themselves (e.g., names, other attendees) to mitigate privacy concerns. Once a user enters the information, the system generates an initial time suggestion that maximizes the availability of the desired attendees. The system returns the time suggestion and initiates a chat so that users can interact with the system to refine the suggested meeting time. For every new chat message entered in the chat, the Constraint Management Component (Section 4.1) translates the request into an action for the system to take. One such action is the addition of a new constraint (see Section 4.2), which requires translating a natural language preference into a Python function. These Python functions are then used within a constraint programming solver (Section 4.3) to generate new time suggestions. An overview of the system is included in Figure 3.
@@
[…]
Three components: User, Chat, and System. The User sends messages via Chat to the System, which returns a response via Chat to the User. The System has four subcomponents: (A) Constraint Manager, (B) System Operations, (C) Prioritized Constraint List, and (D) Constraint Programming Solver. Messages enter the System via the Constraint Manager, feeding into System Operations, which includes Add Constraint, Delete Constraint, Change Priority, Generate Suggestion, and Message User. Add Constraint, Delete Constraint, and Change Priority feed into Prioritized Constraint List, which includes Natural Language Queries mapped to Python Functions. Both Generate Suggestion and Prioritized Constraint List feed into Constraint Programming Solver. Both Message User and Constraint Programming Solver feed into the System’s response back to the User.
@@ 4.1. Constraint Management
  • Add Constraint: This action generates a new scheduling constraint and priority, and calls the constraint generation functionality outlined in Section 4.2.
@@
A key function of the constraint management component is to set the priority of each constraint, and thus maintain the relative ordering of constraints. Upon adding a new constraint, this constraint manager is given access to the current list of scheduling constraints and their priorities and asked to specify the priority of the new constraint. This gives the LLM the flexibility to adjust the priority based on the wording of the new constraint. Based on the chat history, the constraint management component can also decide to change the priority of a constraint (e.g., increasing the probability of satisfying a violated constraint that the user says is important). The system also enables users to change the priority of the constraint during the scheduling process (e.g., if they want a constraint to have higher or lower priority), which helps support preference evolution over time. Figure 5 shows example actions taken by the constraint management component in response to different user messages. A limitation of this priority management functionality is that it remains too opaque to the end user - giving little insight into why a constraint was given a certain priority or access to the prioritized list itself. We further discuss the challenges of using LLMs with respect to system transparency in Section 7.1.1.
@@
[…]
Three input components: Documentation, Examples, and Current Instance as well as an Output component. In this sample, Documentation includes: ‘# Instructions. You are a meeting scheduling assistant managing a list of scheduling constraints. As you talk with the user add and remove constraints to the list. You can also reorder constraints (lower number constraints have higher priority). To change the preference list you can call the following actions: -ADD. This action adds a constraints to the list with the specified priority. The input is a json: ”constraint”: name, ”priority”: priority. The tool returns whether or not the constraint was added successfully, if not it returns why…’ Examples include: ‘Here are some examples: Preference List: [”id”: ”594”, ”priority”: 1, ”preference”: ”Meeting should be scheduled during working hours”, ”id”: ”439”, ”priority”: 2, ”preference”: ”Jennifer should be at the meeting”, ”id”: ”392”, ”priority”: 3, ”preference”: ”Enock should be at the meeting”]. User: The meeting cannot be in the morning. Action: ”ACTION”: ”ADD”, ”INPUT”: ”constraint”: ”No meetings in the morning”, ”priority”: 1. Observation: Constraint added successfully…’ Current Instance includes: ‘# Current Chat History. User: I have to meet before 11am. Action:’ Output includes: ‘”ACTION”: ”ADD”, ”INPUT”: ”constraint”: ”Meeting before 11am”, ”priority”: 1.’ From Current Instance, the line ‘User: I have to meet before 11am’ is highlighted in orange. The entire Output is highlighted in green.
Figure 5. Three sample actions (represented by rows) taken by the constraint management component to manage constraint priorities. In each example, the left image shows the existing prioritized constraint list and current time suggestion. The middle image shows the new user message and the action taken by the constraint management component. The right image shows the resulting prioritized list of constraints and the new time suggestion. In the first example, a user specifies a new constraint with strong language (i.e., ‘need’) that is translated into a constraint with priority 1. The second example shows the same constraint with weaker language (i.e., ‘if possible’) that is translated into a constraint with priority 4. The final example shows a user specifiying that an unment constraint (Anton’s attendance) needs to be met. The constraint manager changes the priority of that constraint to 1.
[…]
Three sample actions (represented by rows) taken by the constraint management component to manage constraint priorities. In each example, the left image shows the existing prioritized constraint list and current time suggestion. The middle image shows the new user message and the action taken by the constraint management component. The final image shows the resulting prioritized list of constraints and the new time suggestion. In the first example, a user specifies a new constraint with strong language (i.e., ‘need’) that is translated into a constraint with priority 1. The second example shows the same constraint with weaker language (i.e., ‘if possible’) that is translated into a constraint with priority 4. The final example shows a user specifiying that an unment constraint (Anton’s attendance) needs to be met. The constraint manager changes the priority of that constraint to 1.
@@ 4.2. Constraint Generation
When the system decides to add a new scheduling constraint, the system calls two LLM-powered components. The first, dubbed Information Checker, checks whether the given scheduling constraint can be handled by the system. Given the diversity of user scheduling preferences and the wide range of external information needed to integrate them, it is unlikely that any system will be able to handle all scheduling constraints. The Information Checker acts as a safeguard to check whether the system has sufficient information to be able to handle the constraint. Within the prompt, the information checker is given all information and data sources accessible to the coding module (e.g., attendee information, function to check each user’s free/busy time) and asked whether the given constraint (e.g., ‘I want to meet before I leave for vacation’) can be coded with an associated rationale (see Figure 6 for a sample prompt). If the information checker predicts that the constraint cannot be implemented with the given data sources, the system generates a new observation for the Constraint Manager component that states that the constraint was not added and gives the information checker module’s rationale as to why. This enables the constraint manager to query additional information needed to implement the constraint (e.g., ask user to specify when they leave for vacation). If the system can handle the constraint, it is then given to a Coder module that translates the scheduling constraint into Python code that can be used to check whether a candidate meeting time meets the constraint. Figure 7 shows a sample prompt for generating a new scheduling constraint.
Figure 6. Sample prompt for the information checker component. Orange highlight reflects query-specific inputs. Green highlights the output of the LLM.
[…]
Three input components: Documentation, Examples, and Current Instance as well as an Output component. In this sample, Documentation includes:You are a meeting scheduling assistant. You have access to the following inputs: - organizer: a string representing the name of the meeting organizer. - attendees: A list of strings representing the attendees of the meeting. …The duration of the meeting is fixed. Your goal is to determine if you have the information needed to code a python function that accomplishes the query. Your answer should be formatted as follows:”response”: yes or no,”rationale”: the reason why you think this is the correct answer.
Figure 7. Sample prompt for the constraint generation coder component. Yellow highlight reflects scheduling instance-specific inputs. Green highlights the output of the LLM.
[…]
Three input components: Documentation, Examples, and Current Instance as well as an Output component. In this sample, Documentation includes: ‘You are a meeting scheduling assistant that is translating a user constraint to code. You have access to the following inputs: -organizer: a string representing the name of the meeting organizer. -duration: an integer representing the duration of the meeting in minutes. -candidate_time: the time to evaluate. Each candidate time has the following: … Your job is to write a python function called meeting_constraint that checks where or not the time meets the condition.’ Examples include: ‘Here are some examples: User: I can only meet in the morning. Code: def meeting_constraint(organizer, duration, candidate_time, calendar_service): return candidate_time.start.hour + duration/60 ¡ 12 …’ Current Instance includes: ‘The meeting you are scheduling has the following details: organizer: Desiree Cain. attendees: Desiree Cain, Collin Lopez, Lauren Sanchez. User: Meeting before 11am. Code:’ Output includes: ‘def meeting_constraint(organizer, duration, candidate_time, calendar_service): return candidate_time.start.hour ¡ 11.’ From Current Instance, the lines organizer: Desiree Cain. attendees: Desiree Cain, Collin Lopez, Lauren Sanchez. User: Meeting before 11am. Code:’ are highlighted in orange. The entire Output is highlighted in green.